OpenStudy (anonymous):

Norman's camping tent has a volume of 72,576 cubic inches. What is the volume of the tent in cubic feet?

OpenStudy (jdoe0001):

\(\large \begin{array}{cclll} 72,576\\ \quad \\ \quad \\ 1ft & = & 12in\\ 1ft^2 & = & 12^2in^2\\ 1ft^3& = & 12^3in^3\\ \quad \\ 72,576 \cancel{in^3}\times \cfrac{ft^3}{12^3\cancel{in^3}} \end{array}\)